# HodgeFormer: Transformers for Learnable Operators on Triangular Meshes through Data-Driven Hodge Matrices

### v4

@@ B.1 Robustness Analysis
To further evaluate the robustness of the HodgeFormer architecture, we perform an analysis, where we evaluate on meshes with added noise, different topology, removed triangles as well as on incomplete meshes. Specifically, we produce variants of the Human dataset’s test set as following:
  • Gaussian Noise: We add Gaussian Noise (GN) $\epsilon$, of several levels $\lambda\in{0.005,0.01,0.02}$ calculated w.r.t. the diagonal of the axis-aligned bounding box of each model, i.e. $\epsilon=\lambda\cdot|BB_{diag}|$.
@@
Incomplete Meshes: Additionally, we select two mesh models from the test set of the Human Body Dataset, one with accurate and one with inaccurate segmentation results and gradually remove body parts from them, similar to [10]. Then, we produce segmentations using a pre-trained HodgeFormer network from Tab.5 as showcased in Fig. S5. The segmentation results remain reasonable and consistent with the initial predictions of the network on these mesh models, even on the one with inaccurate segmentation.
@@ B.4 Variation Study on Label Smoothing
This section examines the effect of label smoothing in the cross-entropy loss on the model’s performance for classification and segmentation tasks. Table S5 reports reports classification results on SHREC11 and segmentation results on COSEG Vases under different label smoothing values.
@@ B.6 Hodgeformer & Transformer Layer Layouts
[…]
Figure S8: HodgeFormer architecture variations. For Human dataset, our highest performing model was a $6$-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, as presented in (a), whereas for Vases, Chairs and Aliens datasets, our highest performing model was a $6$-layered HodgeFormer model followed by two Transformer layers, as presented (b).

### v5

@@ B.1 Robustness Analysis
To further evaluate the robustness of the HodgeFormer architecture, we perform an analysis, where we evaluate on meshes with added noise, different topology, removed triangles as well as on incomplete meshes. Specifically, we produce variants of the Human dataset’s test set as follows:
  • Gaussian Noise: We add Gaussian Noise (GN) $\epsilon$, of several levels $\lambda\in\{0.005,0.01,0.02\}$ calculated w.r.t. the diagonal of the axis-aligned bounding box of each model, i.e. $\epsilon=\lambda\cdot|BB_{diag}|$.
@@
Incomplete Meshes: Additionally, we select two mesh models from the test set of the Human Body Dataset, one with accurate and one with inaccurate segmentation results, and gradually remove body parts from them, similar to [10]. Then, we produce segmentations using a pre-trained HodgeFormer network as showcased in Fig. S5. The segmentation results remain reasonable and consistent with the initial predictions of the network on these mesh models, even on the one with inaccurate segmentation.
@@ B.4 Variation Study on Label Smoothing
This section examines the effect of label smoothing in the cross-entropy loss on the model’s performance for classification and segmentation tasks. Table S5 reports classification results on SHREC11 and segmentation results on COSEG Vases under different label smoothing values.
@@ B.5 Eigenfunctions of the corresponding learned operator
Figure S6: Visualization of mixing strategies of HodgeFormer and vanilla Transformer layers used in Sec. B.3. $N_{H}$ corresponds to the number of HodgeFormer transformer layers and $R_{(N_{H},N_{T})}$ corresponds to the ratio of HodgeFormers to vanilla transformers: a) $N_{H}=4$ & $R_{(N_{H},N_{T})}=4:0$, b) $N_{H}=4$ & $R_{(N_{H},N_{T})}=4:2$, c) $N_{H}=4$ & $R_{(N_{H},N_{T})}=2:1$, d) $N_{H}=4$ & $R_{(N_{H},N_{T})}=1:1$
@@ B.6 Hodgeformer & Transformer Layer Layouts
Figure S8: HodgeFormer architecture variations. For Human dataset, our highest performing model was a $6$-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, as presented in (a), whereas for Vases, Chairs and Aliens datasets, our highest performing model was a $6$-layered HodgeFormer model followed by two Transformer layers, as shown in (b).